# Continuity of asymptotic entropy on wreath products

### v1

@@ 6.2. Increments at unstable points
$\Delta_{n}(b,j)\coloneqq\begin{cases}\varphi_{(j-1)t_{0}}(b)^{-1}\varphi_{jt_{0}}(b),&\text{ if }I_{j}\text{ is an }(R,L)\text{-good interval, and}\\ *,&\text{ otherwise.}\end{cases}$
@@
$\mathbb{E}_{\mu_{k}}\left(\left|\mathcal{V}_{n}\right|\right)<\frac{\varepsilon n}{\log(|L^{t_{0}}|+1)}n.$
@@
Finally, since this inequality holds for all $\varepsilon>0$, we can consider the limit as $\varepsilon$ tends to 00 and obtain $\liminf_{k\to\infty}h(\mu_{k})\geq h(\mu).$ Since the inequality $\limsup_{k\to\infty}h(\mu_{k})\leq h(\mu)$ always holds (see Proposition 3.6), we conclude that
@@ 8.1. Continuity of asymptotic entropy as a consequence of the continuity of harmonic measures
[…]
@@
[…]
@@ 8.2.1. Gromov hyperbolic groups and acylindrically hyperbolic groups
Recall that a geodesic metric space $X$ is called Gromov hyperbolic if there is $\delta\geq 0$ such that every geodesic triangle in $X$ is $\delta$-thin, meaning that any side is contained in a $\delta$-neighborhood of the union of the other two sides. To each hyperbolic space there is an associated Gromov boundary $\partial X$, which consists of equivalence classes of geodesic rays, where two rays are identified if they are at finite Hausdorff distance from each other. On what follows, we will suppose that the space $X$ is separable, which guarantees that the Gromov boundary $\partial X$ carries a natural topology that is separable and completely metrizable. If one furthermore supposes that $X$ is locally compact, then both $\partial X$ and $X\cup\partial X$ are also compact. As a general reference for Gromov hyperbolic spaces, we refer to [BH99, Section III.H.3] and [Väi05]. A finitely generated group is said to be Gromov hyperbolic if one (equivalently any) of its Cayley graphs with respect to a finite generating set is $\delta$-hyperbolic, for some $\delta\geq 0$. For more background on hyperbolic groups we refer to Gromov’s original work [Gro87], and to [GdlH90].
@@
Let us recall that the action of $G$ on $X$ is called acylindrical if for every $K\geq 0$ there are $N,R\geq 1$ such that for any $x,y\in X$ with $d_{X}(x,y)\geq R$, there are at most $N$ elements $g\in G$ such that $d_{X}(x,g.x)\leq K$ and $d_{X}(y,gy)\leq K.$ It is proved in [CFFT22, Theorem 1.2] that, if $\mu$ has finite entropy, then $(\partial X,\lambda)$ is the Poisson boundary of $(G,\mu)$, where $\lambda$ is the hitting measure on $\partial X$. This generalizes [MT18, Theorem 1.5] which had the extra assumption of a finite logarithmic moment of $\mu$, as well as previous descriptions of the Poisson boundary for particular classes of groups that fall within the category of acylindrically hyperbolic groups. In particular, the above covers free groups [DM61, Theorem 3] [Der75, Théorème 2], hyperbolic groups [Anc87, Theorem 8] [Kai94, Theorem 8] [Kai00, Theorems 7.4 & 7.7], groups with infinitely many ends [Woe89, Theorem 7.1] [Kai00, Theorem 8.4], mapping class groups [KM96, Theorem 2.3.1] and $\mathrm{Out}(F_{n})$ [Hor16, Theorem 0.2 & 3.3] [NPR14, Theorem 1.3]. The combination of Theorem 1.5, Proposition 8.6, and the description of the Poisson boundary for acylindrically hyperbolic groups gives us the following result.
@@ 8.2.2. Discrete subgroups of linear groups
Let $G$ be a countable discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$, $d\geq 2$, and consider the associated flag variety $\mathcal{F}$ (see [Led85, Section I.1] for the definition). It is proved in [Led85, Theorem A] [Kai00, Theorem 10.3] that, if $\mu$ is a non-degenerate probability measure on $G$ with a finite first moment, then there is a unique $\mu$-stationary probability measure $\lambda$ on $\mathcal{F}$ such that $(\mathcal{F},\lambda)$ is the Poisson boundary of $(G,\mu)$. A more general version of this statement is proved in [Kai85, Theorem 5] [Kai00, Theorem 10.7] for discrete subgroups semi-simple Lie groups, and which in particular covers the case where $G$ is a Zariski dense discrete subgroup of an algebraic group and $\mu$ is a non-degenerate probability measure on $G$ with finite entropy and a finite logarithmic moment. The above results together with Theorem 1.5 imply the following.
@@
Let $G$ be a countable group. Consider $\mu$, $\mu_{k}$, $k\geq 1$, non-degenerate probability measures on $G$, and such that $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ and $\lim_{k\to\infty}H(\mu_{k})=H(\mu)$. Suppose that we are in one of the following situations.
@@
    $G$ is a countable discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$, $d\geq 2$, and the probability measures $\mu$, $\mu_{k}$, $k\geq 1$, have a finite first moment.
@@
    $G$ is a Zariski dense discrete subgroup of an algebraic group and the probability measures $\mu$, $\mu_{k}$, $k\geq 1$, have finite entropy and a finite logarithmic moment.

### v2

@@ 6.2. Increments at unstable points
[…]
@@
$\mathbb{E}_{\mu_{k}}\left(\left|\mathcal{V}_{n}\right|\right)<\frac{\varepsilon n}{\log(|L^{t_{0}}|+1)}.$
@@
Finally, since this inequality holds for all $\varepsilon>0$, we can consider the limit as $\varepsilon$ tends to 0 and obtain $\liminf_{k\to\infty}h(\mu_{k})\geq h(\mu).$ Since the inequality $\limsup_{k\to\infty}h(\mu_{k})\leq h(\mu)$ always holds (see Proposition 3.6), we conclude that
@@ 8.1. Continuity of asymptotic entropy as a consequence of the continuity of harmonic measures
$\int_{X}f(x)d\nu_{k}=\sum_{g\in G}\mu_{k}(g)\int_{X}f(gx)d\nu_{k}(x).$ (25)
By definition of weak convergence, for the term on the left side of Equation (25), we have $\int_{X}f(x)d\nu_{k}(x)\xrightarrow[k\to\infty]{}\int_{X}f(x)d\nu(x).$ Now let us show that, for the term on the right side of Equation (25), we have $\sum_{g\in G}\mu_{k}(g)\int_{X}f(gx)d\nu_{k}(x)\xrightarrow[k\to\infty]{}\sum_{g\in G}\mu(g)\int_{X}f(gx)d\nu(x).$
@@
$\left|\sum_{g\in G}\mu_{k}(g)\int_{X}f(gx)d\nu_{k}(x)-\sum_{g\in G}\mu(g)\int_{X}f(gx)d\nu(x)\right|$ $\leq\sum_{g\in F}\mu_{k}(g)\left|\int_{X}f(gx)d\nu_{k}(x)-\int_{X}f(gx)d\nu(x)\right|+$
$+\sum_{g\in G\backslash F}\mu_{k}(g)\left|\int_{X}f(gx)d\nu_{k}(x)-\int_{X}f(gx)d\nu(x)\right|+$
$+\sum_{g\in G}\left|\mu_{k}(g)-\mu(g)\right|\left|\int_{X}f(gx)d\nu(x)\right|$
@@ 8.2.1. Gromov hyperbolic groups and acylindrically hyperbolic groups
Recall that a geodesic metric space $X$ is called Gromov hyperbolic if there is $\delta\geq 0$ such that every geodesic triangle in $X$ is $\delta$-thin, meaning that any side is contained in a $\delta$-neighborhood of the union of the other two sides. To each hyperbolic space there is an associated Gromov boundary $\partial X$, which consists of equivalence classes of geodesic rays, where two rays are identified if they are at finite Hausdorff distance from each other. On what follows, we will suppose that the space $X$ is separable, which guarantees that the Gromov boundary $\partial X$ carries a natural topology that is separable and completely metrizable. If one furthermore supposes that $X$ is locally compact, then both $\partial X$ and $X\cup\partial X$ are also compact. As a general reference for Gromov hyperbolic spaces, we refer to [BH99, Section III.H.3] and [VÄI05]. A finitely generated group is said to be Gromov hyperbolic if one (equivalently any) of its Cayley graphs with respect to a finite generating set is $\delta$-hyperbolic, for some $\delta\geq 0$. For more background on hyperbolic groups we refer to Gromov’s original work [GRO87], and to [Gd90].
@@
Let us recall that the action of $G$ on $X$ is called acylindrical if for every $K\geq 0$ there are $N,R\geq 1$ such that for any $x,y\in X$ with $d_{X}(x,y)\geq R$, there are at most $N$ elements $g\in G$ such that $d_{X}(x,g.x)\leq K$ and $d_{X}(y,gy)\leq K.$ It is proved in [CFF+25, Theorem 1.2] that, if $\mu$ has finite entropy, then $(\partial X,\lambda)$ is the Poisson boundary of $(G,\mu)$, where $\lambda$ is the hitting measure on $\partial X$. This generalizes [MT18, Theorem 1.5] which had the extra assumption of a finite logarithmic moment of $\mu$, as well as previous descriptions of the Poisson boundary for particular classes of groups that fall within the category of acylindrically hyperbolic groups. In particular, the above covers free groups [DM61, Theorem 3] [DER75, Théorème 2], hyperbolic groups [ANC87, Theorem 8] [KAI94, Theorem 8] [KAI00, Theorems 7.4 & 7.7], groups with infinitely many ends [WOE89, Theorem 7.1] [KAI00, Theorem 8.4], mapping class groups [KM96, Theorem 2.3.1] and $\mathrm{Out}(F_{n})$ [HOR16, Theorem 0.2 & 3.3] [NPR14, Theorem 1.3]. The combination of Theorem 1.5, Proposition 8.6, and the description of the Poisson boundary for acylindrically hyperbolic groups gives us the following result.
@@ 8.2.2. Discrete subgroups of linear groups
Let $G$ be a countable discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$, $d\geq 2$, and consider the associated flag variety $\mathcal{F}$ (see [LED85, Section I.1] for the definition). It is proved in [LED85, Theorem A] [KAI00, Theorem 10.3] that, if $\mu$ is a non-degenerate probability measure on $G$ with a finite first moment, then there is a unique $\mu$-stationary probability measure $\lambda$ on $\mathcal{F}$ such that $(\mathcal{F},\lambda)$ is the Poisson boundary of $(G,\mu)$. A more general version of this statement is proved in [KAI85, Theorem 5] [KAI00, Theorem 10.7] for discrete subgroups semi-simple Lie groups, and which in particular covers the case where $G$ is a Zariski dense discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$ and $\mu$ is a non-degenerate probability measure on $G$ with finite entropy and a finite logarithmic moment. The above results together with Theorem 1.5 imply the following.
@@
Let $G$ be a countable discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$, $d\geq 2$. Consider $\mu$, $\mu_{k}$, $k\geq 1$, non-degenerate probability measures on $G$, and such that $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ and $\lim_{k\to\infty}H(\mu_{k})=H(\mu)$. Suppose that we are in one of the following situations.
@@
    The probability measures $\mu$, $\mu_{k}$, $k\geq 1$, have a finite first moment.
@@
    The group $G$ is Zariski dense and the probability measures $\mu$, $\mu_{k}$, $k\geq 1$, have finite entropy and a finite logarithmic moment.
@@ 8.2.5. Free-by-cyclic groups
8.2.6. Wreath products
One can also apply Theorem 1.5 to recover a particular case of Theorem 1.2 regarding the continuity of asymptotic entropy on wreath products. We first prove the next convergence of harmonic measures on lamplighter boundaries.
Proposition 8.12.
Let $A$, $B$ be countable groups and let $\mu,\mu_{k}$, $k\geq 1$, be non-degenerate probability measures on $A\wr B$ that define a transient random walk on $B$. Suppose that
  1. (1)
    $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ for each $g\in G$,
  2. (2)
    there is a finite subset $F\subseteq B$ such that $\mathrm{supp}(\mu_{k})\subseteq\{(f,b)\in A\wr B\mid\mathrm{supp}(f)\subseteq F\}$ for each $k\geq 1$, and
  3. (3)
    $\lim_{k\to\infty}p_{\mathrm{esc}}(\pi_{*}\mu_{k})=p_{\mathrm{esc}}(\pi_{*}\mu)$, where $\pi:A\wr B\to B$ is the canonical projection to the base group $B$.
Denote by $\nu$, $\nu_{k}$, $k\geq 1$, their corresponding hitting measures on $\prod_{B}A$. Then $\nu_{k}\xrightarrow[k\to\infty]{}\nu$ weakly.
Item (2) above can be interpreted as requiring the sequence of probability measures to have a “uniformly bounded lamp range”, meaning that there exists a fixed finite neighborhood in which the lamplighter person may modify lamps around their current position at each step. Note that this condition together with the transience assumption for the random walk in the base group guarantees the almost sure stabilization of lamp configurations, so that the hitting measures $\nu$, $\nu_{k}$, $k\geq 1$, are well defined. Note also that in this case the space $\prod_{B}A$ will in general not be uniquely $\mu$-stationary. As an example, one can think of the case where $A$ is finite and $B$ is amenable. Then $\prod_{B}A$ is a compact space that must possess an $A\wr B$-invariant probability measure, due to the amenability of $A\wr B$, which will be different from the hitting measure of a random walk.
Proof of Proposition 8.12.
The space $\prod_{B}A$ has a countable basis for its topology consisting of unions of cylinders of the form
$\mathcal{C}_{\mathbf{p}}\coloneqq\left\{f:B\to A\mid f(g)=\mathbf{p}(g)\text{ for each }g\in\mathrm{dom}(\mathbf{p})\right\},$
where $\mathbf{p}:\mathrm{dom}(\mathbf{p})\to A$ is a function from a finite subset $\mathrm{dom}(\mathbf{p})\subseteq B$ to $A$.
To show that $\nu_{k}\xrightarrow[k\to\infty]{}\nu$ weakly, it suffices to prove that for any such cylinder as above we have $\nu_{k}(\mathcal{C}_{\mathbf{p}})\xrightarrow[k\to\infty]{}\nu(\mathcal{C}_{\mathbf{p}}).$ Indeed, this follows from the fact that any open set in $\prod_{B}A$ can be written as a countable union of disjoint cylinders, together with the Portmanteau theorem [BIL99, Theorem 2.1].
Recall that we denote by $(\varphi_{n},X_{n})$, $n\geq 1$, a trajectory of a random walk on $A\wr B$. Let us introduce the event
$E_{\infty}=\{\varphi_{\infty}(g)=\mathbf{p}(g)\text{ for each }g\in\mathrm{dom}(\mathbf{p})\},$
and for each $N\geq 1$ the events
$E_{N}=\{\varphi_{N}(g)=\mathbf{p}(g)\text{ for each }g\in\mathrm{dom}(\mathbf{p})\},$
$T_{N}=\{X_{n}\notin\mathrm{dom}(\mathbf{p})\cdot F^{-1}\text{ for all }n\geq N\},$
where $F\subseteq B$ is a finite subset such that $\mathrm{supp}(\mu_{k})\subseteq\{(f,b)\in A\wr B\mid\mathrm{supp}(f)\subseteq F\}$ for each $k\geq 1$. This is a hypothesis in Item (2) above.
Then for each $N\geq 1$ we have
$\nu_{k}(\mathcal{C}_{\mathbf{p}})=\mathbb{P}_{\mu_{k}}(E_{\infty})=\mathbb{P}_{\mu_{k}}(E_{N}\cap T_{N})+\mathbb{P}_{\mu_{k}}(E_{\infty}\cap T_{N}^{c}).$
Note also that we have
$\mathbb{P}_{\mu_{k}}(E_{N}\cap T_{N})=\sum_{b\in B}\mathbb{P}_{\mu_{k}}\left(E_{N}\cap\{X_{N}=b\}\cap\{X_{N}^{-1}X_{n}\notin b^{-1}\mathrm{dom}(\mathbf{p})F^{-1}\text{ for all }n\geq N\}\right),$
and conditioning together with the independence of the increments, we get
$\mathbb{P}_{\mu_{k}}(E_{N}\cap T_{N})=\sum_{b\in B}\mathbb{P}_{\mu_{k}}\left(E_{N}\cap\{X_{N}=b\}\right)\mathbb{P}_{\mu_{k}}\left(\{X_{n}\notin b^{-1}\mathrm{dom}(\mathbf{p})F^{-1}\text{ for all }n\geq 0\}\right).$
Let $\varepsilon>0$. First, let us choose $N\geq 1$ sufficiently large such that $\mathbb{P}_{\mu_{k}}(T_{N}^{c})<\varepsilon$ for each $k\geq 1$ and $\mathbb{P}_{\mu}(T_{N}^{c})<\varepsilon$. The existence of such an $N$ is guaranteed by the continuity of the escape probability of the random walk in the base group, assured by Item (3) in the statement above. Next, we can find a finite subset $Z\subseteq B$ and $K_{1}\geq 1$ have $\mathbb{P}_{\mu_{k}}(X_{N}\notin Z)<\varepsilon$ for all $k\geq K_{1}$ and $\mathbb{P}_{\mu}(X_{N}\notin Z)<\varepsilon$. This is assured by the convergence of Item (1) above (see also Lemma 3.1). Using again the convergence of escape probabilities in the base group, and using the convergence from Item (1) and Lemma 3.2 which guarantees the convergence of the convolutions of the measures, we can find $K_{2}\geq K_{1}$ such that for each $k\geq K_{2}$ we have
$\left|\begin{aligned} &\sum_{b\in Z}\mathbb{P}_{\mu_{k}}\left(E_{N}\cap\{X_{N}=b\}\right)\mathbb{P}_{\mu_{k}}\left(\{X_{n}\notin b^{-1}\mathrm{dom}(\mathbf{p})F^{-1}\text{ for all }n\geq 0\}\right)-\\ &-\sum_{b\in Z}\mathbb{P}_{\mu}\left(E_{N}\cap\{X_{N}=b\}\right)\mathbb{P}_{\mu}\left(\{X_{n}\notin b^{-1}\mathrm{dom}(\mathbf{p})F^{-1}\text{ for all }n\geq 0\}\right)\end{aligned}\right|<\varepsilon$
By combining the above equations and using the triangular inequality several times, we conclude that
$\left|\nu_{k}(\mathcal{C}_{\mathbf{p}})-\nu(\mathcal{C}_{\mathbf{p}})\right|\leq 5\varepsilon$
for each $k\geq K_{2}$. Since $\varepsilon>0$ was arbitrary, this shows the desired weak convergence. ∎
Let $A$ be a countable group and let $B$ be a hyper-FC-central countable group. It was proved in [FS23, Corollary 1.4] that for any probability measure $\mu$ on $A\wr B$ with $H(\mu)<\infty$ and such that the lamp configurations stabilize almost surely, the space $\prod_{B}A$ endowed with the hitting measure is the Poisson boundary of $(A\wr B,\mu)$. This result generalized previous work that identified the Poisson boundary of $A\wr B$ under stronger assumptions both on the measure $\mu$ as well as on the groups $A$ and $B$ by A. Erschler [ERS11b, Theorem 1] and R. Lyons and Y. Peres [LP21, Theorem 1.1]. Combining this with Proposition 8.12 and with Theorems 1.4 and 1.5 we obtain the following particular case of Theorem 7.1.
Corollary 8.13.
Let $A$ be a countable group and let $B$ be a hyper-FC-central countable group with at least cubic growth. Let $\mu,\mu_{k}$, $k\geq 1$, be non-degenerate probability measures on $A\wr B$ with finite entropy. Suppose that
  1. (1)
    $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ for each $g\in G$,
  2. (2)
    $\lim_{k\to\infty}H(\mu_{k})=H(\mu)$,
  3. (3)
    there is a finite subset $F\subseteq B$ such that $\mathrm{supp}(\mu_{k})\subseteq\{(f,b)\in A\wr B\mid\mathrm{supp}(f)\subseteq F\}$ for each $k\geq 1$.
Then $\lim_{k\to\infty}h(\mu_{k})=h(\mu)$.
We remark that this approach cannot recover Theorem 1.2 in full generality for arbitrary sequences of probability measures with finite entropy. At this level of generality, there is no identification of the Poisson boundary of $A\wr B$, let alone one that simultaneously identifies the Poisson boundaries associated with different probability measures within a common Polish model space. In particular, the hitting measure on the space $\prod_{B}A$ is not well defined, since lamp configurations need not stabilize in this level of generality; see the last paragraph of Subsection 2.3.